OpenStudy (anonymous):

Can someone please tell me how to solve this word problem? :) Two cars leave Orlando heading towards Atlanta.The first car is going at 50 mph and the second car is going 70 mph. The first car leaves at 9AM and the second car leaves at 11AM. How long will it take for the second car to pass the first car?

OpenStudy (anonymous):

The second car will pass the first one when it will have covered the same distance as the first car. Although the times will of course vary. Let X be the distance covered by car A (the first one) and car B (the second one). Time taken by A to cover X = Distance / Speed = X / 50 Time taken by B to cover X = X/70 + 2hours [--2] (It started at 11 am remember? 2 hours after the first one) Now equating both the times, X/50 = X/70 5 2 X/50 = (X + 140)/70 {I just took the LCM} 70X = 50 (x+140) {Cross multiplying} 70X = 50X + 7000 20X = 7000 X = 350. Therefore the car B will pass car A after covering 350 miles. Hence time taken will be, (putting value of X in eq. --2) = 350/70 + 2 = 5 +2 = 7 hours :)

OpenStudy (anonymous):

NOW THATS A DAM GOOD ANSWER !! WOOOOO :D Thats what I got!

OpenStudy (anonymous):

Haha :D

OpenStudy (anonymous):

Wahhhahahahahaha >:)

OpenStudy (anonymous):

In this problem equate the distances traveled from the start line. r t = d Let a clock start when the second car begins to move. At this point in time the first car is 100 miles away, 50*2. Solve the following for t: 70 t = 100 + 50 t t = 5 hours The second car will intercept the first car at 4PM.

OpenStudy (anonymous):

I got 7 hours and so did my mom and ifcmaniac?

OpenStudy (anonymous):

You messed up with the times Robtobey. the second car takes t + 2 hours. Because it started 2 hours later. :)

OpenStudy (anonymous):

ARe you sure?

OpenStudy (anonymous):

Yes. 9AM + 7 hours is 1600 hours or 4PM 11AM + 5 hours is 1600 hours or 4PM

OpenStudy (anonymous):

Both cars travel 350 miles, the second car travels 5 hours @ 70 mph and the second travels 7 hours @ 50 mph.